Fractions equivalent to 4/10

Fractions equivalent to 4/10?

are 1/2 and 4/10 equivalent fractions?
Answer #1
Fractions equivalent to 4/10 =

2/5
8/20
12/30
16/40

Do you see the pattern?

4/10 >> Divide both the numerator and denominator by 2 to get 2/5

4/10 >> Multiply the numerator and denomiator by 2 to get 8/20
answered by: Monti
Answer #2
No they are not.
1/2 = 5/10
Since 4/10 ≠ 5/10, they are not equivalent fractions.
answered by: glynn